Theory and Applications of Categories, Vol. 9

PARTICIPANTS IN THE CT2000 CONFERENCE

Jirì Adamek (Technical University Braunschweig) Steven Awodey (Carnegie Mellon University) Michael Batanin (Macquarie University) Andrej Bauer (Carnegie Mellon University) Tibor Beke (Universiteit Utrecht) Lars Birkedal (The IT University of Copenhagen) Steven Bloom (Stevens Institute) Richard Blute (University of Ottawa) Francis Borceux (Université Catholique de Louvain) Dominique Bourn (Université du Littoral) Ronald Brown (University of Wales) Stefaan Caenepeel (Vrije Universiteit Brussel) Aurelio Carboni (Università dell’Insubria) Claudia Centazzo (Dalhousie University) Eugenia Cheng (University of Cambridge) Maurizio Citterio (Università di Milano) Robin Cockett (University of Calgary) Bob Coecke (Vrije Universiteit Brussel) Sjoerd Crans (McGill University) Tamar Datuashvili (Georgian Academy of Science) Robert Dawson (St. Mary’s University) Federico De Marchi (University of Leicester) Francisco Javier Diaz (Universidad de la Laguna) Yves Diers (Université de Valenciennes) Eduardo Dubuc (University of Buenos Aires) John Duskin (SUNY/Buffalo) Jeff Egger (University of Cambridge) Zoltán Ésik (A. Jozsef University, Szeged) Peter Freyd (University of Pennsylvania) Jonathon Funk (University of Saskatchewan) Jose Manuel Garcia Calcines (Universidad de la Laguna) Dale Garraway (Dalhousie University) Robbie Gates (Macquarie University) Marek Golasinski (Uniwersytet M. Kopernika)

Received by the editors 2002 February 21.

Published on 2002 March 10 in the volume of articles from CT2000.

\begin{matrix}c \\ ◯ \end{matrix} The Editors of Theory and Applications of Categories, 2001. Permission to copy for private use

granted.

Marco Grandis (Università di Genova) Marino Gran (Université Catholique de Louvain) John Gray (University of Illinois) Michael Healy (The Boeing Company) Alex Heller (CUNY) Claudio Hermida (IST, Universidade Tecnica Lisboa) Anne Heyworth (University of Leicester) Peter Hines (University of Wales) Martin Hyland (University of Cambridge) George Janelidze (Georgian Academy of Science) Michael Johnson (Macquarie University) Peter Johnstone (University of Cambridge) Heiner Kamps (Fern Universität Hagen) Panagis Karazeris (University of Patras) Christian Kassel (Institut de Recherche Mathématique Avancée, Strasbourg) Piergiulio Katis (Università dell’Insubria) G. Max Kelly (University of Sydney) Rudger Kieboom (Vrije Universiteit Brussel) Heinrich Kleisli (University of Fribourg) Anders Kock (Aarhus Universitet) Anna Labella (Università di Roma) Stephen Lack (University of Sydney) F. William Lawvere (SUNY/Buffalo) Tom Leinster (University of Cambridge) Fred Linton (Wesleyan University) Saunders Mac Lane (University of Chicago) John MacDonald (University of British Columbia) Maria Emilia Maietti (Università di Padova) Florence Marty (Université Paris VII) Luca Mauri (Université de Montpellier) Paddy McCrudden (Macquarie University) Matías Menni (University of Edinburgh) Marie-Anne Moens (Université Catholique de Louvain) Ieke Moerdijk (Universiteit Utrecht) Susan Niefield (Union College) Zaza Omiadze (Georgian Academy of Science) Robert Paré (Dalhousie University) Cristina Pedicchio (Università di Trieste) Alberto Peruzzi (Università di Firenze) Markus Pfenniger (IpMP) Timothy Porter (University of Wales) Pedro Resende (IST, Universidade Tecnica Lisboa)

Robert Rosebrugh (Mount Allison University) Jiri Rosický (Masarykova Univerzita) Giuseppe Rosolini (Università di Genova) George Rousseau (University of Leicester) Luigi Santocanale (Aarhus Universitet) Robert Seely (McGill University) Haim Shapira (College of Management, Rishon Le Zion) Alex Simpson (University of Edinburgh) Pawel Sobocinski (University of Sydney) Steven Sourbron (Vrije Universiteit Brussel) Lurdes Sousa (Istituto Politecnico de Viseu) Richard Squire (McGill University) Daniel Steffen (Macquarie University) Thomas Streicher (TU Darmstadt) Isar Stubbe (Université Catholique de Louvain) Paul Taylor (Queen Mary and Westfield College) Michel Thiebaud (College de Stael) Walter Tholen (York University) Todd Trimble (Loyola University Chicago) Enrico Vitale (Université Catholique de Louvain) Rainer Vogt (Universität Osnabrück) Robert Walters (University of Sydney/Università dell’Insubria) Mark Weber (Macquarie University) Ralph Wojtowicz (Rose-Hulman Institute of Technology) Michael Wright Noson Yanofsky (CUNY) Marek Zawadowski (Uniwersytet Warszawski) Jaap van Oosten (Universiteit Utrecht)

This list may be accessed via WWW at http://www.tac.mta.ca/tac/ or by anonymous

ftp at ftp://ftp.tac.mta.ca/pub/tac/html/volumes/9/list/list.{dvi,ps}

THEORY AND APPLICATIONS OF CATEGORIES (ISSN 1201-561X) will disseminate articles that significantly advance the study of categorical algebra or methods, or that make significant new contributions to mathematical science using categorical methods. The scope of the journal includes: all areas of pure category theory, including higher dimensional categories; applications of category theory to algebra, geometry and topology and other areas of mathematics; applications of category theory to computer science, physics and other mathematical sciences; contributions to scientific knowledge that make use of categorical methods.

Articles appearing in the journal have been carefully and critically refereed under the responsibility of members of the Editorial Board. Only papers judged to be both significant and excellent are accepted for publication.

The method of distribution of the journal is via the Internet tools WWW/ftp. The journal is archived electronically and in printed paper format.

Subscription information. Individual subscribers receive (by e-mail) abstracts of articles as they are published. Full text of published articles is available in .dvi, Postscript and PDF. Details will be e-mailed to new subscribers. To subscribe, send e-mail to tac@mta.ca including a full name and postal address. For institutional subscription, send enquiries to the Managing Editor, Robert Rosebrugh,

rrosebrugh@mta.ca.

Information for authors. The typesetting language of the journal is TEX, and L^{A}TEX is the preferred flavour. TEX source of articles for publication should be submitted by e-mail directly to an appropriate Editor. They are listed below. Please obtain detailed information on submission format and style files from the journal’s WWW server at http://www.tac.mta.ca/tac/. You may also write to tac@mta.ca to receive details by e-mail.

Editorial board.

John Baez, University of California, Riverside: baez@math.ucr.edu Michael Barr, McGill University: barr@barrs.org, Associate Managing Editor Lawrence Breen, Université Paris 13: breen@math.univ-paris13.fr Ronald Brown, University of North Wales: r.brown@bangor.ac.uk Jean-Luc Brylinski, Pennsylvania State University: jlb@math.psu.edu Aurelio Carboni, Università dell Insubria: aurelio.carboni@uninsubria.it P. T. Johnstone, University of Cambridge: ptj@dpmms.cam.ac.uk G. Max Kelly, University of Sydney: maxk@maths.usyd.edu.au Anders Kock, University of Aarhus: kock@imf.au.dk F. William Lawvere, State University of New York at Buffalo: wlawvere@acsu.buffalo.edu Jean-Louis Loday, Université de Strasbourg: loday@math.u-strasbg.fr Ieke Moerdijk, University of Utrecht: moerdijk@math.uu.nl Susan Niefield, Union College: niefiels@union.edu Robert Paré, Dalhousie University: pare@mathstat.dal.ca Andrew Pitts, University of Cambridge: Andrew.Pitts@cl.cam.ac.uk Robert Rosebrugh, Mount Allison University: rrosebrugh@mta.ca, Managing Editor Jiri Rosicky, Masaryk University: rosicky@math.muni.cz James Stasheff, University of North Carolina: jds@math.unc.edu Ross Street, Macquarie University: street@math.mq.edu.au Walter Tholen, York University: tholen@mathstat.yorku.ca Myles Tierney, Rutgers University: tierney@math.rutgers.edu Robert F. C. Walters, University of Insubria: walters@fis.unico.it R. J. Wood, Dalhousie University: rjwood@mathstat.dal.ca